User:Puja Mody/Notebook/Chem 571: Gold Nanoparticles/2012/10/17

From OpenWetWare
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

Gel Electrophoresis


  • Run the plasmids through a gel electrophoresis.
  • Complete HRP chemiluminescence assay


  • Two solutions were created. In each one, 1μL of DPN1 was added to 45μL K110A (f) and K110A (r) plasmid respectively. These plasmid solutions were the same ones made and placed in the thermocycler Yesterday. The two new solutions were then placed on a heat block at 37°C for 60 minutes.
  • Next, 5μL of the two solutions K110A (f) and K110A (r) plasmids were combined with 1μL of 6X Gel Loading Dye (blue). the 6μL solutions were then loaded into an agarose gel for gel electrophoresis.
  • Gel Electrophoresis
    • 1.23 grams of agarose was mixed with 35mL of 1X TAE buffer (buffer made of Tris base, Acetic acid and Ethylenediaminetetraacetic acid (EDTA) and then microwaved for 45 seconds. This solution was then quickly poured into the Gel electrophoresis and was given approximately 20-30 minutes to solidify. 1X TAE buffer was then poured into the gel electrophoresis container.
    • A DNA ladder was made by mixing 5uL DNA ladder with 1uL of 6X gel loading dye blue. The Plasmids with different ADA mutations were then loaded in the gel. There are a total of 3 different plasmids each one with a f (forward) and an r (reverse). Therefore, between the entire lab, there were 6 plasmids. They were loaded in gel in the following order:
      • DNA ladder, E34K, E34A , E34A, K110A, K110A, E34K
    • The Gel electrophoresis was run at 85 volts for 1 hour.

As I was not present for the next segment of the experiment, these steps and information was taken from Keyun Wang's Notebook

  • "Gel was stained in TAE buffer with ethidium bromide. The gel, as well as 20uL of 12.3mg/mL ethidium bromide were added into 100mL of TAE buffer and mixed for 5 minutes. Gel was then transferred to TAE buffer without ethidium bromide to rinse for 5 minutes."
  • "Gel was observed under UV lamp to see if PCR product has been successfully produced."


  • The Fluorimeter was not working today and attempts to use the Uv-Vis in a makeshift manner in order to calculate the luminescence emitted were unsuccessful. This was done by blocking the source lamp so that only the emitted light could be measured. However, data collection by this method was not possible as no signal was detected. As a result this experiment has been placed on hold until the equipment is once again working.
  • The duration of lab was used to make new Au/BSA solutions at the same original 14 concentrations and the lysozyme stock solution was also made. The lysozyme will be used to nucleate the AuNPs in the future. I was not present in lab during this time. For information regarding the making of the Au/BSA solutions please consult Michaels Notebook. For instructions and calculations regarding the lysozyme stock please look at Mary's Notebook.